# DAViD: Modeling Dynamic Affordance of 3D Objects using Pre-trained Video Diffusion Models

### v1

@@ 3 Method
To obtain the remaining object motion, we first transform the default object pose with a transformation that transforms the camera used for object rendering to match the estimated camera $\Pi_{d}^{f_{1}}$ in 3D space as follows:
$\bm{R}_{o^{\dagger}}^{f_{1}}=\bm{R}_{\text{cam},d}^{f_{1}}\left(\bm{R}_{\gamma(d)}\right)^{-1}\bm{R}_{o}$ (2)
$\bm{t}_{o^{\dagger}}^{f_{1}}=\bm{R}_{\text{cam},d}^{f_{1}}\left(\bm{R}_{\gamma(d)}\right)^{-1}(\bm{t}_{o}-\bm{t}_{\gamma(d)})+\bm{t}_{\text{cam},d}^{f_{1}},$ (3)
where $\gamma(\cdot)$ is the function that maps the index of the 2D HOI image to the index of the camera from which it is rendered. $R_{o^{\dagger}}^{f_{1}},t_{o^{\dagger}}^{f_{1}}$ are each of the transformed object rotation and translation which is aligned with the first frame of the generated 2D HOI video. This is possible only for the first frame since the relative pose between camera and object used for object rendering corresponds only to the generated 2D HOI image, which is the first frame of the 2D HOI video. Our intuition for determining the object’s pose for later frames is to find the corresponding points of the object among 2D video frames, obtaining cues for the 3D pose of the object. We first identify the 3D object vertices $v_{o,\text{3D}}\in\mathbb{R}^{n_{o}\times 3}$ that are visible from the camera in the first frame $\Pi_{d}^{f_{1}}$ by using raycasting [74]. For corresponding 2D rendered points $\bm{v}_{o,\text{2D}}^{f_{1}}=\Pi_{d}^{f_{1}}(\bm{v}_{o,\text{3D}})\in\mathbb{R}^{n_{o}\times 2}$ of the visible 3D object vertices $\bm{v}_{o,\text{3D}}$ , we utilize a 2D correspondence estimator [37, 36] to obtain the 2D corresponding points $\{\bm{v}_{o,\text{2D}}^{f_{i}}\}_{i=2}^{F}$ in the remaining video frames. As we have the 3D object vertices and the corresponding 2D matching points for each frame, we compute the camera poses for each frame with respect to the fixed object pose $\{\bm{R}_{o^{\dagger}}^{f_{1}},\bm{t}_{o^{\dagger}}^{f_{1}}\}$ using Perspective-n-Points (PnP) [20, 47] as follows:
$\{\bm{R}_{\text{rel}}^{f_{i}},\bm{t}_{\text{rel}}^{f_{i}}\}=\text{PnP}(\bm{v}_{o,\text{2D}}^{f_{i}},\bm{v}_{o,\text{3D}}).$ (4)
Using the estimated camera pose $\Pi_{d}^{f_{i}}$ in world coordinate, we obtain the final object poses for each frame, $\{R_{o^{\dagger}}^{f_{i}},t_{o^{\dagger}}^{f_{i}}\}_{i=1}^{F}$ using the camera pose computed via PnP [20, 47] and applying a reverse transformation as follows.
$\bm{R}_{o^{\dagger}}^{f_{i}}=\bm{R}_{\text{cam},d}^{f_{i}}\left(\bm{R}_{\text{rel}}^{f_{i}}\right)^{-1}\bm{R}_{o^{\dagger}}^{f_{1}}$ (5)
$\bm{t}_{o^{\dagger}}^{f_{i}}=\bm{R}_{\text{cam},d}^{f_{i}}\left(\bm{R}_{\text{rel}}^{f_{i}}\right)^{-1}(\bm{t}_{o^{\dagger}}^{f_{1}}-\bm{t}_{\text{rel}})+\bm{t}_{\text{cam},d}^{f_{i}}.$ (6)
Resolving Scale-Depth Ambiguity. Although we uplift human motion and object motion into a common 3D space, they are positioned along the rays so that their projections are matched to the 2D rendering, without considering their actual depth locations due to the scale-depth ambiguity. To resolve the ambiguity, we optimize the scale of the human ($s_{h}$) and object ($s_{o}$) in the camera direction using (1) a weak depth cue from the depth map, and (2) a contact cue from the object movement assumption. The loss term of weak depth cue is defined as follows:
$\mathcal{L}_{h}$ $=\left\lVert\left(s_{h}(\bm{v}_{h,\text{3D}}-\bm{t}_{h,d}^{f_{1}})+\bm{t}_{h,d}^{f_{1}}\right)-\bm{P}_{h,\text{3D}}\right\rVert_{2}^{2}$ (7)
$\mathcal{L}_{o}$ $=\left\lVert\left(s_{o}(\bm{v}_{o,\text{3D}}-\bm{t}_{o,d}^{f_{1}})+\bm{t}_{o,d}^{f_{1}}\right)-\bm{P}_{o,\text{3D}}\right\rVert_{2}^{2},$ (8)
where $\bm{v}_{h,\text{2D}}\in\mathbb{R}^{n_{h}\times 2}$ , $\bm{v}_{h,\text{3D}}\in\mathbb{R}^{n_{h}\times 3}$ are visible 2D and 3D vertices of SMPL-X [69] in $\mathcal{I}_{d}$ computed similarly as $\bm{v}_{o,\text{2D}}$, $\bm{v}_{o,\text{3D}}$, and $\bm{P}_{h,\text{3D}}$, $\bm{P}_{o,\text{3D}}$ are 3D point cloud generated by projecting 2D metric depths (computed by metric depth model [6]) of $\bm{v}_{h,\text{2D}}$ $\bm{v}_{o,\text{2D}}$ into 3D space. Note that $\bm{t}_{h,d}^{f_{1}}\in\mathbb{R}^{n_{h}\times 3}$, $\bm{t}_{o,d}^{f_{1}}\in\mathbb{R}^{n_{o}\times 3}$ are expanded matrices of $\bm{t}_{\text{cam},d}^{f_{1}}\in\mathbb{R}^{3}$, where each row of $\bm{t}_{h,d}^{f_{1}},\bm{t}_{o,d}^{f_{1}}$ is the same as $\bm{t}_{\text{cam},d}^{f_{1}}$. The loss term of the contact cue is defined as follows:
$\mathcal{L}_{\text{HOI}}=d_{n}(\bm{P}_{H},\bm{P}_{O})$ (9)
$\bm{P}_{H}=s_{h}(\bm{V}_{H}-\bm{t}_{H,d}^{f_{1}})+\bm{t}_{H,d}^{f_{1}}$ (10)
$\bm{P}_{O}=s_{o}(\bm{V}_{O}-\bm{t}_{O,d}^{f_{1}})+\bm{t}_{O,d}^{f_{1}},$ (11)
where $d_{n}(\bm{P}_{x},\bm{P}_{y})$ is the average distance of the $n$ closest points of $\bm{P}_{y}$ to $\bm{P}_{x}$, and $\bm{V}_{O}\in\mathbb{R}^{N_{o}\times 3}$, $\bm{V}_{H}\in\mathbb{R}^{N_{h}\times 3}$ are point clouds consist of every vertex of object and human. Note that $n$ is a hyperparameter that varies across categories, and $\bm{t}_{H,d}^{f_{1}}\in\mathbb{R}^{N_{h}\times 3}$, $\bm{t}_{O,d}^{f_{1}}\in\mathbb{R}^{N_{o}\times 3}$ are also expanded matrices of $\bm{t}_{\text{cam},d}^{f_{1}}\in\mathbb{R}^{3}$ where each row of $\bm{t}_{H,d}^{f_{1}},\bm{t}_{O,d}^{f_{1}}$ is the same as $\bm{t}_{\text{cam},d}^{f_{1}}$. Assuming that the object requires contact with the human to have movement, we set the loss to encourage the human and object to have contact points. We optimize the scale $s_{h}$, $s_{o}$ to obtain $s_{h}^{*}$, $s_{o}^{*}$ by minimizing total loss as follows:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{h}+\mathcal{L}_{o}+\mathcal{L}_{\text{HOI}}$ (12)
As the original scale of human motion output by GVHMR [81] is already matched to real-world scale, we keep the scale of the human fixed and adjust the scale of the object by a factor of $s_{o}^{*}/s_{h}^{*}$. Since we use a perspective camera, the depth is also adjusted by the same ratio.
@@ 3.2 Learning Dynamic Affordance
Given the diverse dynamic 4D HOI samples we synthesized from a target 3D object, we pursue to build a generative motion model to learn the dynamic patterns of human and the target object. To learn such affordance, we propose DAViD, which consists of (1) LoRA [33] module to leverage a pre-trained MDM [88], and (2) human-conditioned object pose diffusion model. Our DAViD models both the distribution of human motion, and the distribution of object poses conditioned on human poses, enabling to generate diverse 4D HOI motion samples within the distributions.
Figure 4: Qualitative Results. DAViD models the dynamic patterns of both a human and an object for various object categories. We visualize HOI motions for various objects generated by our DAViD, where frames are visualized in temporal order. The results show that our method effectively models the dynamic patterns emerging in HOIs.
LoRA Module for MDM. To learn the patterns of human motion interacting with a specific object, we insert a LoRA [33] layer into a pre-trained MDM [88] model and train the layer for each object. By choosing LoRA [33], we learn human motion patterns that are unseen in the pre-trained model, even with a small amount of training data. More interestingly, we find that by merging the trained LoRA [33] layers, it is possible to combine multiple human motion patterns during HOIs with a single model, which is previously unseen in the data samples.
In practice, we first consider the human motion part only without the object parts from 4D HOI samples previously generated, and preprocess the motion data following the method of HumanML3D [25], obtaining the training data of human joints denoted as $\bm{j}\in\mathbb{R}^{F\times 22\times 3}$. Then, the LoRA [33] layer is added to the multi-head attention layer of the transformer encoder block in text-conditioned MDM [88] as shown in Fig. 3, and trained with the following joint recovery loss function.
@@
where $\phi$ is the original weight of pre-trained MDM [88] denoted as $\mathcal{M}$, and $\Delta\phi$ is the weight of our LoRA [33] module we want to obtain. For the training text prompt, we use automatically simplified text compared to the original used in 2D HOI image generation. See supplementary material for the detailed process.
Human Conditioned Object Pose Diffusion. Once we synthesize a plausible human motion, we then add a corresponding 3D object motion. As humans are the driving force behind an object’s movement, the motion of the object is inherently conditioned on the human’s motion. Specifically, we observe that the human pose at a specific timeframe is closely correlated to the object pose at the same time instance. Based on this intuition, we first (1) train a diffusion model that generates object poses conditioned on human poses, and then (2) provide smoothness guidance at inference time for the human motion condition, which we named Smoothness Guidance Sampling (SGS). In practice, we first sample $1024$ vertices from the T-pose of SMPL-X [69] using Poisson disk sampling [58]. For the sampled indices, we normalize each frame of the generated 4D HOI sample by adjusting the translation so that the mean of the sampled human vertices aligns with the origin, and rotate the root orientation to align with z-axis of the world coordinate. Note that this transformation is applied to both the human and the object. A pair of the SMPL-X [69] body pose parameters $\bm{\theta}\in\mathbb{R}^{21\times 3}$ and the normalized object pose $\bm{p}\in\mathbb{R}^{3\times 4}$ (consists of rotation and translation) is used for training our diffusion model. Using the dataset, our objective is to learn model $\Psi$ which outputs the score function of conditional object pose distribution for all noise level $t$.
$\Psi(\bm{p},t|\bm{\theta})=\nabla_{\bm{p}}\log p(\bm{p}_{t}|\bm{\theta})$ (14)
According to Pascal et al. [94], the objective can be achieved by minimizing the following loss function.
$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(\epsilon,1)}\lambda(t)\mathbb{E}_{\bm{p},t}\left\lVert\Psi_{\theta}(\bm{p}_{t},t|\bm{\theta})-\frac{\bm{p}_{0}-\bm{p}_{t}}{\sigma(t)^{2}}\right\rVert_{2}^{2}$ (15)
@@
where $\sigma_{\text{min}}=0.1$, $\sigma_{\text{max}}=50$ are hyperparameters about noise perturbation. For the detailed structure of our score model, we follow the architecture of GenPose [113]. After training the conditional object pose diffusion model, we propose a simple but powerful technique to generate motion-like trajectories of the object at inference time called Smoothness Guidance Sampling (SGS). We estimate the object pose sequence $\bm{p}^{f_{1}:f_{F}}\in\mathbb{R}^{F\times 3\times 4}$, which is conditioned on the body pose sequence $\bm{\theta}^{f_{1}:f_{F}}\in\mathbb{R}^{F\times 21\times 3}$. Instead of being guided solely on the estimated score from the model, we add a smoothness guidance term into the Probability Flow (PF) ODE [83] as follows:
$\begin{gathered}\frac{d\bm{p}^{f_{1}:f_{F}}}{dt}=-\sigma(t)\dot{\sigma}(t)\Psi(\bm{p}^{f_{1}:f_{F}},t,\bm{\theta}{f_{1}:f_{F}})\\ +\text{SGS}(\bm{p}^{f_{1}:f_{F}}_{t}),\end{gathered}$ (17)
where SGS($\bm{p}^{f_{1}:f_{F}}_{t}$) is the gradient of recovered $\hat{\bm{p}}_{0}^{f_{1}:f_{F}}\in\mathbb{R}^{F\times 3\times 4}$ from the given noise level of object pose on following smoothness loss.
$\mathcal{L}_{s}(\hat{\bm{p}}_{0}^{f_{1}:f_{F}})=\left\lVert\hat{\bm{p}}_{0}^{f_{2}:f_{F}}-\hat{\bm{p}}_{0}^{f_{1}:f_{F-1}}\right\rVert_{2}^{2}$ (18)
Note that we utilize RK45 ODE solver [17] for figuring out the ODE trajectory. SGS allows to find temporally smooth object pose trajectories while sampling each frame-wise pose in trained distribution.
@@ 4 Experiments
In this section, we conduct experiments to evaluate the efficacy of our method. In Section 4.3, we show various generated dynamic 3D HOI samples, demonstrating the efficacy of DAViD on learning dynamic patterns, and in Section 4.4, we conduct a quantitative comparison with other baselines. Finally, in Section 4.5, we introduce an application which shows the potential of our pipeline.
@@ 4.1 Datasets
For qualitative evaluation, we consider 30 input 3D object meshes from various sources, obtained from BEHAVE [5], InterCap [34], ShapeNet [10, 60], SAPIEN [105], FullBodyManipulation Dataset [49], and SketchFab [2]. For quantitative evaluation, we use FullBodyManipulation Dataset [49], which consists of the motion of both humans and objects. As our goal is to learn dynamic concepts of HOI using a few shots of 4D HOI samples, we conduct evaluations on object categories with relatively fewer long-tail interactions. Specifically, we use the mop object for the evaluation.

### v2

@@ 3 Method
Intuitively, the human and object at time $f$ are observed from the same camera, and thus we can simply co-locate them in the same space, by transferring each component to its camera coordinate, where the camera becomes the origin:
$\bm{\hat{V}}_{o}^{f}=\bm{R}_{o}^{f}\bm{V}_{o}^{f}+\bm{t}_{o}^{f},$ (3)
$\bm{\hat{V}}_{h}^{f}=\bm{R}_{h}^{f}\bm{V}_{h}^{f}+\bm{t}_{h}^{f},$ (4)
where $\bm{V}_{h}^{f}\in\mathbb{R}^{3}$ is a human 3D vertex in its global coordinate after applying all SMPL parameters, and $\bm{\hat{V}}_{o}^{f}$ and $\bm{\hat{V}}_{h}^{f}$ are the human and object 3D vertices in the shared camera coordinate. To this end, we place both the 3D human and the 3D object in the global space of the human side (defined by GVHMR) to obtain their global 4D movements.
Resolving Depth-Scale Ambiguity. Although we uplift and place human and object movements in a common 3D space, there still exists the depth-scale ambiguity, since each components are defined in arbitrary object scales, as shown in depth-scale optimization part in Fig. 3. To resolve the ambiguity, we optimize the scale of the human ($s_{h}\in\mathbb{R}$) and object ($s_{o}\in\mathbb{R}$) in the camera coordinate at the first frame, using (1) a weak depth cue from the depth map, and (2) a contact cue from the object movement assumption. The objective function is defined as follows:
$\mathcal{L}_{\text{total}}=\lambda_{h}\mathcal{L}_{h}+\lambda_{o}\mathcal{L}_{o}+\lambda_{\text{HOI}}\mathcal{L}_{\text{HOI}}+\lambda_{\text{col}}\mathcal{L}_{\text{collision}},$ (5)
where $\lambda$s denote the weighting factors. The $\mathcal{L}_{h}$ and $\mathcal{L}_{o}$ are the depth consistency loss for human and object, defined as:
$\mathcal{L}_{h}$ $=\sum_{i\in\text{vis}_{\text{h}}}\left\lVert s_{h}\bm{\hat{V}}_{h,i}-\bm{P}_{h,i}\right\rVert^{2}$ (6)
$\mathcal{L}_{o}$ $=\sum_{j\in\text{vis}_{\text{o}}}\left\lVert s_{o}\bm{\hat{V}}_{o,j}-\bm{P}_{o,j}\right\rVert^{2},$ (7)
where $\bm{\hat{V}}_{h,i}$ and $\bm{\hat{V}}_{o,i}$ are the $i$-th human and object vertices in the camera coordinate at the first frame respectively, and $\bm{P}_{h,\text{3D}}$, $\bm{P}_{o,\text{3D}}$ are the corresponding 3D points by back-projecting 2D metric depths (computed by metric depth model [4]) into the camera coordinate. Note that the index $i,j$ are from the set of visible vertex indices of human and object at the first frame.
Additionally, the loss term of the weak HOI contact cue is defined as follows:
$\mathcal{L}_{\text{HOI}}=\mathcal{D}_{n}\left(\{s_{h}\bm{\hat{V}}_{h,i}\},\{s_{o}\bm{\hat{V}}_{o,j}\}\right),$ (8)
where $\mathcal{D}_{n}$ is the function to compute the average distance of the $n$ closest points between two points sets. Intuitively, the cost term encourages contact between the human and object, assuming they are in contact at the first frame. Note that $n$ is a hyperparameter that varies across categories. We apply the collision term to avoid undesired penetration using COAP [59]. The object vertices are used as query points, reducing the collision between human and object in $\mathcal{L}_{\text{collision}}$.
@@ 3.2 Learning Dynamic Affordance
Based on the diverse dynamic 4D HOI samples we synthesized for a target 3D object, we build a generative HOI motion model, DAViD, to model dynamic affordance. Our DAViD consists of two key components: (1) a generative human motion model by fine-tuning a pre-trained MDM [88] with object-specific LoRA [31] module, and (2) a score-based diffusion model for estimating object poses conditioned by generated human poses. For the score-based diffusion model, we introduce a novel inference-time sampling technique, TGS to enhance overall quality of motion.
Figure 4: Qualitative Results. DAViD models the dynamic patterns of both a human and an object for various object categories. We visualize HOI motions for various objects generated by our DAViD, where frames are visualized in temporal order.
Human Motion Synthesis with LoRA for MDM. To capture human motion patterns interacting with a specific object, we insert a LoRA [31] layer into a pre-trained MDM [88] model and train it for each object category. By leveraging LoRA [31], our model can learn previously unseen human motion patterns with a limited amount of training samples, keeping the original motion synthesis quality of the pre-trained model. Interestingly, we find that merging LoRA [31] layers trained accross different object categories enables our model to generate human motion interacting with multiple objects in a zero-shot manner (see Fig. 5 (c)).
For training, we extract only the human motion from our 4D HOI samples, excluding object motions. We preprocess the motion data following HumanML3D [24], and obtain the training data of human joints denoted as $\bm{j}\in\mathbb{R}^{F\times 22\times 3}$. The LoRA [31] layer is then added into the multi-head attention layer of the transformer encoder block in text-conditioned MDM [88] as shown in Fig. 3. Our model is trained with the following joint recovery loss function.
@@
where $\phi$ is the original weight of pre-trained MDM [88] denoted as $\mathcal{M}$, $c$ is the encoded text prompt using CLIP encoder [73], and $\Delta\phi$ is the weight of our LoRA [31] module we want to obtain. For the training text prompt, we use the category name and annotate the additional hand index (left or right) for categories with hand-object interactions to improve controllability. See Supp. Mat. for more details.
Object Motion Synthesis from Human Motion. After generating plausible human motion for a given 3D object, we synthesize a corresponding 3D object motion conditioned on the human motion. To achieve it, we design our object motion module with two key parts: (1) a score-based diffusion model for modeling the distribution of object pose conditioned on the corresponding 3D human pose in HOI, (2) an inference method to estimate the final motion trajectory across a sequence of frames, with a novel sampling technique, named Temporal Guidance Sampling (TGS), for enhancing motion consistency and fine contact details.
To train the score-based diffusion model, we first normalize each frame of the generated 4D HOI sample by adjusting the translation so that the pelvis joint aligns with the origin, and rotate the root orientation to align with z-axis of the world coordinate. The transformations are applied to both the human and the object. A pair of the SMPL-X [69] body pose parameters $\bm{\theta}\in\mathbb{R}^{21\times 3}$ and the normalized object pose $\bm{p}\in\mathbb{R}^{9}$ (consists of 6D rotation and 3D translation) are used for training our diffusion model. Using the pre-processed data samples, we train our score-based diffusion model $\Psi$ to output the score of the conditional object pose ($\bm{p}_{t}$) distribution for all timestep $t$:
$\Psi(\bm{p},t|\bm{\theta})=\nabla_{\bm{p}}\log p(\bm{p}_{t}|\bm{\theta}).$ (10)
As demonstrated in Pascal et al. [94], the objective can be achieved by minimizing the following loss function.
$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(\epsilon,1)}\lambda(t)\mathbb{E}_{\bm{p},t}\left\lVert\Psi(\bm{p}_{t},t|\bm{\theta})-\frac{\bm{p}_{0}-\bm{p}_{t}}{\sigma(t)^{2}}\right\rVert_{2}^{2}$ (11)
@@
where $\sigma_{\text{min}}=0.1$, $\sigma_{\text{max}}=50$ are hyperparameters about noise perturbation. For the detailed structure of our score model, we follow the architecture of GenPose [114].
Based on our conditional object pose diffusion model, we present an inference method to estimate the entire object motion sequence, rather than sampling object poses independently per frame. This process is further enhanced by our Temporal Guidance Sampling (TGS) technique, which encourages temporal consistency and human-object contact during inference. Specifically, we estimate the object pose sequence over F frames $\bm{p}^{1:F}\in\mathbb{R}^{F\times 9}$, conditioned on the corresponding body motion sequence $\bm{\theta}^{1:F}\in\mathbb{R}^{F\times 21\times 3}$. Instead of relying solely on the estimated score from the model, we introduce additional guidance term into the Probability Flow ODE [84] as follows:
$\begin{gathered}\frac{d\bm{p}^{1:F}}{dt}=-\sigma(t)\dot{\sigma}(t)\Psi(\bm{p}^{1:F},t|\bm{\theta}{1:F})+\nabla_{\bm{p}^{1:F}}{\mathcal{L_{\text{TGS}}}}.\end{gathered}$ (13)
The $\nabla_{\bm{p}^{1:F}}{\mathcal{L_{\text{TGS}}}}$ is the gradient of the object pose at time step $t$ on the following TGS loss, $\mathcal{L}_{\text{TGS}}$:
$\mathcal{L}_{\text{TGS}}(\bm{p}^{1:F})=\lambda_{\text{s}}\mathcal{L}_{\text{smooth}}+\lambda_{\text{c}}\mathcal{L}_{\text{contact}},$ (14)
$\mathcal{L}_{\text{smooth}}(\bm{p}^{1:F})=\left\lVert\bm{p}^{2:F}-\bm{p}^{1:{F-1}}\right\rVert_{2}^{2},$ (15)
$\mathcal{L}_{\text{contact}}(\bm{p}^{1:F})=\mathcal{D}_{\text{chamfer}}\left(\mathcal{V}_{h},\mathcal{V}_{o}(\bm{p}^{1:F})\right),$ (16)
where $\mathcal{L}_{\text{smooth}}$ encourages temporal consistency between consecutive frames. The $\mathcal{L}_{\text{contact}}$ encourages proximity between human and object vertices, measured by the Chamfer Distance $\mathcal{D}_{\text{chamfer}}$ between the human vertices $\mathcal{V}_{h}$ and object vertices $\mathcal{V}_{o}$ positioned by the current pose $\bm{p}^{1:F}$. In practice, we use a certain threshold to focus on nearby points only when computing the Chamfer Distance. We utilize RK45 ODE solver [16] to solve the ODE trajectory.
@@ 4 Experiments
In this section, we conduct experiments to evaluate our method. In Section 4.3, we show 4D HOI motions generated by DAViD for various objects, demonstrating the efficacy of DAViD in learning dynamic HOI patterns. In Section 4.4, we conduct a quantitative comparison with other baselines and a quantitative ablation study on our design choices.
@@ 4.1 Datasets
For qualitative evaluation, we consider 30 input 3D object meshes from various sources, obtained from BEHAVE [3], InterCap [32], ShapeNet [8, 60], SAPIEN [106], FullBodyManipulation Dataset [50], and SketchFab [83]. For quantitative evaluation, we use FullBodyManipulation Dataset [50], which consists of the motion of both humans and objects.